40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4164     Entry time: Mon Jan 17 20:13:38 2011
Author: rana 
Type: Update 
Category: Electronics 
Subject: POX_11 Optical TF 

I used 50 mA to drive the laser diode. The light is split 50/50 between the DUT (Device Under Test) and the New Focus 1611 (1 GHz BW) diode used as the reference.

This measurement is the TF of DUT/(New Focus). The resonances are there, but clearly there's an issue with instability around 200 MHz. The setup is still powered up, so please be careful around the RFPD testing table (don't stomp around yank the cables out of the power supplies).

I looked at the RF Photodiode wiki that Alberto has started - most of the TF features are replicated there. Todo:

* Update the 'schematic' with a real schematic instead of the cartoon.
* Change the circuit to remove the resistor in the RF path.
* Add compensation to avoid the 200 MHz instability.
* Make sure to include opamp current noise in the noise model (it is the dominant noise source but has been left out in the noise estimation plot).
* Make the output into a true 50 Ohms.
Attachment 1: A.TIF  22 kB  | Hide | Hide all
A.TIF
ELOG V3.1.3-